2. How many pounds of cream containing 12% butterfat must be added to 1800 lbs. of milk containing 2% butterfat to obtain milk having 3% butterfat?

Answer by mananth(16949)

How many pounds of cream containing 12% butterfat must be added to 1800 lbs. of milk containing 2% butterfat to obtain milk having 3% butterfat?
let 12% butter fat added be x (12/100=0.12)
0.12x+1800*0.02 = 0.03(x+1800)
0.12x +36 = 0.03x+54
0.09x=18
x=18/0.09
x=200 lbs